# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ Abstract.
Brownian Boost is a one-parameter family of stochastic differential games played on the real line in which players spend at rates of their choosing in an ongoing effort to influence the drift of a randomly diffusing point particle $X$. One or other player is rewarded, at time infinity, according to whether $X$ tends to plus or minus infinity. Each player’s net receipt is the final reward (only for the victor) minus the player’s total spend. We characterise and explicitly compute the time-homogeneous Markov-perfect Nash equilibria of Brownian Boost, finding the derivatives of the players’ expected payoffs to solve a pair of coupled first-order non-linear ODE. Brownian Boost is a high-noise limit of a two-dimensional family of player-funded tug-of-war games, one of which was studied in [26]. We analyse the discrete games, finding them, and Brownian Boost, to exemplify key features studied in the economics literature of tug-of-war initiated by [27]: a battlefield region where players spend heavily; stakes that decay rapidly but asymmetrically in distance to the battlefield; and an effect of discouragement that makes equilibria fragile under asymmetric perturbation of incentive. Tug-of-war has a parallel mathematical literature derived from [41], which solved the scaled fair-coin game in a Euclidean domain via the infinity Laplacian PDE. By offering an analytic solution to Brownian Boost, a game that models strategic interaction and resource allocation, we seek to build a bridge between the two tug-of-war literatures.
@@
1. Introduction
1.1. Brownian Boost
@@
We begin by introducing the stochastic differential game at the heart of this study, signposting some of the principal inferences we will reach, and stating the analytic framework—an ODE pair with explicit solutions—by which Nash equilibria in Brownian Boost will be classified.
@@ 1.1.1. Game setup
Fix $\rho\in(0,\infty)$. Mina and Maxine play $\rho$-Brownian Boost or $\mathrm{BB}(\rho)$.
@@
${\rm d}X_{t}=\frac{a(t)^{\rho}-b(t)^{\rho}}{a(t)^{\rho}+b(t)^{\rho}}\,{\rm d}t+\,{\rm d}W_{t}\,$,
where $W:[0,\infty)\to\mathbb{R}$ is standard Brownian motion, and $X(0)=0$.
Brownian Boost ends only at time infinity. It does so with left escape if $E_{-}:=\big\{X(t)\to-\infty\big\}$ occurs and with right escape if $E_{+}:=\big\{X(t)\to\infty\big\}$ does; the limits are in high $t$. The respective events represent victory in the game for Mina and Maxine. Maxine receives a terminal receipt $T_{+}:={\bf 1}_{E_{+}}$; Mina, one of $T_{-}:=\lambda\cdot{\bf 1}_{E_{-}}$. That is, a prize goes to the winning player, with none to her opponent (and note that, should escape $E:=E_{-}\cup E_{+}$ fail to occur, then no prize is offered); currency has been revalued so Maxine’s prize is one unit when awarded, and this leaves one parameter, Mina’s victory reward $\lambda\in(0,\infty)$, that completes the specification of the game data.
@@
Players pay for their stakes. Maxine’s running cost is $R_{+}=\int_{0}^{\infty}a(t){\rm d}t$; Mina’s, $R_{-}=\int_{0}^{\infty}b(t){\rm d}t$. These costs are deducted from terminal receipts, so that Maxine and Mina’s net total payoffs from playing the game are equal to $P_{+}:=T_{+}-R_{+}$ and $P_{-}:=T_{-}-R_{-}$.
@@
Suppose that players may choose from a reasonable class of stake strategies adapted to the game’s history until the present moment. Each seeks to maximize her expected net receipt, $\mathbb{E}\,P_{-}$ or $\mathbb{E}\,P_{+}$. For given $\rho\in(0,\infty)$, we may ask for which values of Mina’s terminal reward $\lambda\in(0,\infty)$ do there exist Nash equilibria, or strategy pairs from which neither player would benefit by deviating unilaterally. And if equilibria exist, are they unique, and may they be explicitly described?
The translation-invariant real line may appear to be a featureless terrain on which to seek to secure local geographic advantage, and the dispensing of costly resources in the short term a profligate choice in a game of infinite duration. Yet our study is animated by the presence and structure of Nash equilibria in Brownian Boost, via the themes of the competition between securing territorial advantage and the financial burden incurred in the attempt, and of how the relative incentive of players, manifest in the value of $\lambda$ relative to one, influences the judgements that they make.

### v4

@@ Abstract.
Brownian Boost is a one-parameter family of stochastic differential games played on the real line in which two players continuously expend resources in an effort to influence the drift of a randomly diffusing point particle $X$, with victory determined only at time infinity. Each player chooses a spending rate over time, incurring cumulative cost, and receives a terminal reward if the particle ultimately escapes in her preferred direction. We characterise and explicitly compute the time-homogeneous Markov-perfect Nash equilibria of Brownian Boost, showing that the derivatives of the players’ equilibrium payoff functions solve a pair of coupled first-order non-linear ODE.
Brownian Boost arises as a fine-mesh, high-noise limit of a two-parameter family of discrete, player-funded tug-of-war games, one instance of which was studied in [24]. We analyse the discrete games and Brownian Boost, finding them to exemplify key features studied in the economics literature of tug-of-war initiated by [25]: a battlefield region where players spend heavily; spending rates that decay rapidly but asymmetrically in distance to the battlefield; and an effect of discouragement that renders equilibria fragile under asymmetric perturbation of incentive.
Tug-of-war has a parallel mathematical literature derived from [43], which solved the scaled fair-coin game in a Euclidean domain via the infinity Laplacian PDE. By offering an analytic solution to Brownian Boost—a game modelling strategic interaction and resource allocation—we aim to draw the attention of probabilists and analysts to mathematically novel tug-of-war games whose delicate analytic structure admits economically meaningful interpretation.
@@
Preface
Fix $\rho\in(0,\infty)$. Mina and Maxine play $\rho$-Brownian Boost. A point-particle counter $X:[0,\infty)\to\mathbb{R}$ starts at the origin. Left to its own devices, $X$ is Brownian; but it is equipped with a motor that may impute a drift, left or right, of magnitude at most one. At any given time $t\geq 0$, Maxine and Mina stake money at respective non-negative rates that for convenience we denote by $a(t)$ and $b(t)$, though in principle players’ decisions may depend on the entire counter history until the present time. The counter evolves according to the stochastic differential equation
${\rm d}X_{t}=\frac{a(t)^{\rho}-b(t)^{\rho}}{a(t)^{\rho}+b(t)^{\rho}}\,{\rm d}t+\,{\rm d}W_{t}\,$, (1)
where $W:[0,\infty)\to\mathbb{R}$ is standard Brownian motion (and $X(0)=0$). That is to say, the instantaneously imputed drift equals $\frac{a(t)^{\rho}-b(t)^{\rho}}{a(t)^{\rho}+b(t)^{\rho}}\in[-1,1]$, with Mina or Maxine securing drift closer to 1-1 or +1+1 by dispersing funds more rapidly, and thus dominating more easily if $\rho$ is higher.
The victor of Brownian Boost is decided only at time infinity. It is Mina when $X(t)\to$-$\infty$ and Maxine when $X(t)\to\infty$. The victor alone receives a terminal reward: one for Maxine and $\lambda\in(0.\infty)$ for Mina, so that the given pair $(\rho,\lambda)\in(0,\infty)^{2}$ specifies the game data.
Players pay for their stakes, each accruing a running cost given by the integral over all positive time of her chosen stake rate. Each player’s reserves are great enough as to permit unfettered access to funds, but the accrued cost for a given player is deducted from her (possibly zero) terminal reward. In this way, stake decisions are governed not by spending limits but by the cost of expenditure relative to terminal reward.
This stochastic differential game is a cousin of tug-of-war games that have been the object of study in a long strand of economics research in dynamic contest theory originating with Harris and Vickers [25]. Two reactions may seem natural to the reader.
  • Is the game not unplayable? Brownian Boost is an unusual game from an economic standpoint, because no reward is offered according to how the counter travels over any finite period of time. The choice is natural, forcing any structure in equilibrium to arise endogenously from strategic interaction, rather than from an indelicate preference for some region enforced by design. Yet the emergence of such structure is hardly guaranteed. The real line is a featureless terrain on which to seek to obtain a local geographic advantage by irrevocably dispensing a valued resource. Why would a player spend at all? A player who has outbid her opponent for a long time will likely have induced near-uniform motion at speed close to one, carrying the counter far from the origin in her preferred direction. But what benefit compensates her spent funds when the present location is indistinguishable from any other, according to the translation invariance of the real line? Surely the funds would have better gone unspent.
  • Discouragement for the less incentivized. If $\lambda>1$, then Mina has a greater incentive to win. She may plausibly meet any given stake rate offered by Maxine by proportionally outstaking her, even by a slight margin. The counter would then drift in her favour at a uniform rate, assuring her victory. Maxine would recognize this, accept her certain defeat, and save her running cost; Mina could then win with a very small cost. It would seem then that any asymmetry in incentive is incompatible with non-trivial structure.
As we will see, the first objection is false. The conclusion reached by the second is correct for Brownian Boost but false for counterpart discrete games, though in a weakened guise it retains some accuracy.
A Tullock contest [47] with exponent $\rho\in(0,\infty)$ is a single-round competition in which a player who stakes $a\in(0,\infty)$ against the opponent’s $b\in(0,\infty)$ wins with probability $\frac{a^{\rho}}{a^{\rho}+b^{\rho}}$. A natural discrete stake-governed tug-of-war game may be played on a finite integer interval: each round is a Tullock contest that results in a movement of the counter one step to the left or the right, until one or other end of the gameboard is reached. This game is a close cousin of Harris and Vickers’, in which two firms invest resources in research over time in an ongoing competition to secure a patent.
In this article, we study a family of such tug-of-war games, on a gameboard given by the set of integers rather than on a finite interval therein. These Trail of Lost Pennies games serve as natural discrete counterparts to Brownian Boost. The family has two parameters: first, the Tullock exponent $\rho$; second, a noise parameter $\kappa\in(0,1)$ that governs the strength of mixing of stake-based moves with independent $\pm 1$ moves decided by a fair coin flip.
These games offer a means to rigorously interpret and study Brownian Boost, and they provide a cohesive framework in which to examine natural hypotheses such as the two mooted above, for both the continuum model and the discrete counterparts.
We exhibit Brownian Boost, for $\rho\in(0,1]$, as a fine-mesh, high-noise scaling limit of the discrete games. In this representation, we obtain a classification of the Markov perfect Nash equilibria of Brownian Boost, describing them explicitly in terms of a coupled ODE pair. In the discrete games, we find a rich collection of behaviour at Nash equilibrium, exhibiting robustness under perturbations of incentive from equality that, while positive, is highly fragile. This behaviour illustrates and quantifies economically significant effects such as discouragement, and the presence of battlefield regions wherein players commit resources intensively and away from which these commitments decay rapidly but asymmetrically.
Chapter 1 Introduction
This article has two principal aims: to rigorously analyse Brownian Boost, and to develop and solve a broad class of stake-governed tug-of-war games whose equilibria illuminate themes in dynamic contest theory such as the discouragement effect and robustness (or fragility) under asymmetric perturbations of terminal incentives.
The two aims are pursued in tandem. We introduce a two-parameter family of discrete tug-of-war games on $\mathbb{Z}$, the Trail of Lost Pennies, one instance of which was studied in [24]. A natural solution concept is Markov-perfect Nash equilibrium, under which players seek to maximise mean net receipts, acting solely on the basis of present counter location. We will classify explicitly the resulting equilibria in these discrete games, over a broad swathe of parameter space.
Brownian Boost is then exhibited as a fine-mesh, high-noise scaling limit of the Trail of Lost Pennies. This representation provides a rigorous framework for the stochastic differential game and circumvents the difficulties associated with instantaneous feedback, while the accompanying scaling of the discrete game solutions yields an explicit ODE-based characterization of Brownian Boost equilibria. Together, the discrete models and their continuum limit form a fairly broad menagerie of explicitly solved models with which to analyse themes in dynamic contest theory.
1.1. Brownian Boost: setup and signposts
@@
We develop our presentation of the stochastic differential game at the heart of this study, and signpost one of the principal inferences we will reach, concerning gameplay under the Brownian Boost equilibria classified in the article.
@@ 1.1.1. Game setup
The game ends at time infinity with left escape if $E_{-}:=\big\{X(t)\to-\infty\big\}$ occurs and with right escape if $E_{+}:=\big\{X(t)\to\infty\big\}$ does; the limits are in high $t$. Each player receives a terminal payment according to which event occurs. In principle, four parameters are needed, and we may write Maxine’s terminal receipt $T_{+}$ in the form $m_{\infty}{\bf 1}_{E_{+}}+m_{-\infty}{\bf 1}_{E_{-}}$ and Mina’s $T_{-}$ as $n_{\infty}{\bf 1}_{E_{+}}+n_{-\infty}{\bf 1}_{E_{-}}$. On the vector $(m_{-\infty},m_{\infty},n_{-\infty},n_{\infty})\in\mathbb{R}^{4}$ we impose $m_{-\infty}<m_{\infty}$ and $n_{\infty}<n_{-\infty}$ in order that $E_{-}$ and $E_{+}$ may rightly be viewed as events of victory for Mina and Maxine. In practice, three trivial symmetries serve to reduce this boundary-data parameter space to a one-dimensional subspace. Suitable pre-game payments to the players permit us to set the losing receipts $m_{-\infty}$ and $n_{\infty}$ equal to zero. And, since the drift in (1) is homogeneous of degree zero in the stake pair $(a,b)$, currency revaluation then permits Maxine’s victory receipt to equal one. The remaining free parameter is Mina’s winning receipt, which we set equal to the given value $\lambda\in(0,\infty)$. To wit, we take $T_{+}={\bf 1}_{E_{+}}$ and $T_{-}=\lambda\cdot{\bf 1}_{E_{-}}$, as in the preface.
@@
Maxine’s running cost is $R_{+}=\int_{0}^{\infty}a(t){\rm d}t$; Mina’s, $R_{-}=\int_{0}^{\infty}b(t){\rm d}t$. Maxine and Mina’s net total payoffs from playing the game are thus equal to $P_{+}:=T_{+}-R_{+}$ and $P_{-}:=T_{-}-R_{-}$.
@@
Each player seeks to maximize her expected net receipt, $\mathbb{E}\,P_{-}$ or $\mathbb{E}\,P_{+}$. Each chooses a strategy that determines her stake-rate $a(t)$ or $b(t)$ at every time $t\geq 0$ in a manner adapted to the game’s history until the present moment. A Nash equilibrium is a strategy-pair from which neither player would unilaterally choose to deviate given the declared aim of maximizing mean receipt.
This description of admissible strategies is imprecise, and we will shortly indicate more about the framework in which we interpret and analyse Brownian Boost. For now, it is useful to explain a basic aspect which is needed to explain our analysis of the game. In playing $\mathrm{BB}(\rho)$, a player may in principle draw on a broad range of strategies determined by game history in choosing her stakes. We will restrict attention to a narrower class that includes all viable options according to an intuitively appealing principle akin to the Markov property: in the history of gameplay until a given moment, the one piece of data that should be determinative for deciding the stake rate is the present counter position $X_{t}$. Stake pairs that meet this condition are time-homogeneous and Markov perfect, and in shorthand we will call them time-invariant. By focussing on such pairs, we reinterpret the stakes specified in the preface as profiles $a,b:\mathbb{R}\to[0,\infty)$, with $a(x)$ and $b(x)$ denoting the rate at which Maxine and Mina stake at any moment $t\geq 0$ for which $X_{t}=x$. In this way, the Brownian Boost equilibria that we seek to classify are the time-invariant Nash equilibria (or TINE). Each TINE is a map $(a,b):\mathbb{R}\to[0,\infty)$ whose domain is the gameboard $\mathbb{R}$. A given time-invariant strategy pair, including any TINE, induces gameplay satisfying the SDE (1) recorded in the preface where the drift term takes the form $\frac{a(X_{t})^{\rho}-b(X_{t})^{\rho}}{a(X_{t})^{\rho}+b(X_{t})^{\rho}}$.
Moscarini and Smith [39] study a model whose specification is close to the game we consider; formally, it is obtained from Brownian Boost by setting $\rho=1/2$ and omitting the denominator in the drift term in (1). For games played on finite intervals, the symmetric Markov-perfect equilibria are characterized via Hamilton–Jacobi–Bellman arguments. We will discuss [39] and its relation to our work in Section 5.1 after giving a heuristic derivation of the HJB equations for our system.